Would someone please help me with this problem...
Given f(x)=x^2+3 and g(x)(sqrt)2-x find f o g(x).
a.5-x
b.1-x
c (sqrt)5-x^2
d. 7+x^2

Thanks

Answer by stanbon(75887)

Given f(x)=x^2+3 and g(x)=(sqrt)2-x find f o g(x).
--------
f[g(x)] = f[ sqrt2 - x]
= (sqrt2 - x)^2 + 3
=2-2sqrt2x+x^2 + 3
= x^2 - 2sqrt2x +5
--------
Sorry, this is the answer as you have posted the problem,
but the answer is not among the ones you listed.
----------
Cheers,
Stan H.